### v1

# Frequency-Assisted Mamba for Remote Sensing Image Super-Resolution

Abstract
Recent progress in remote sensing image (RSI) super-resolution (SR) has exhibited remarkable performance using deep neural networks, e.g., Convolutional Neural Networks and Transformers. However, existing SR methods often suffer from either a limited receptive field or quadratic computational overhead, resulting in sub-optimal global representation and unacceptable computational costs in large-scale RSI. To alleviate these issues, we develop the first attempt to integrate the Vision State Space Model (Mamba) for RSI-SR, which specializes in processing large-scale RSI by capturing long-range dependency with linear complexity. To achieve better SR reconstruction, building upon Mamba, we devise a Frequency-assisted Mamba framework, dubbed FMSR, to explore the spatial and frequent correlations. In particular, our FMSR features a multi-level fusion architecture equipped with the Frequency Selection Module (FSM), Vision State Space Module (VSSM), and Hybrid Gate Module (HGM) to grasp their merits for effective spatial-frequency fusion. Recognizing that global and local dependencies are complementary and both beneficial for SR, we further recalibrate these multi-level features for accurate feature fusion via learnable scaling adaptors. Extensive experiments on AID, DOTA, and DIOR benchmarks demonstrate that our FMSR outperforms state-of-the-art Transformer-based methods HAT-L in terms of PSNR by 0.11 dB on average, while consuming only 28.05% and 19.08% of its memory consumption and complexity, respectively.
Index Terms:
Remote sensing image, super-resolution, state space model, frequency selection
I Introduction
High-resolution remote sensing imagery (RSI), which records high-quality earth observation details, provides promising prospects for large-scale and fine-grained applications [1, 2, 3, 4, 5]. However, the complex imaging environment (e.g., scattering and tremor) often impedes high-resolution (HR) image acquisition [6, 7, 8]. Moreover, RSI often undergoes severe compression and downsampling to tame the transmission instability between satellites and ground stations, resulting in suboptimal scene representation. Hence, reconstructing HR images from low-resolution (LR) observations is crucial for improving both human perception and subsequent applications.
In contrast to upgrading hardware maintenance, super-resolution (SR) techniques provide a flexible and cost-effective alternative by predicting latent HR images from their LR counterparts [9, 10, 11, 12]. Early efforts often relied on hand-crafted priors to tame the ill-posedness [13, 14, 15, 16]. However, they struggled to produce accurate results and involved laborious optimization processes. Recently, deep neural networks have demonstrated remarkable progress in SR tasks and achieved superior performance over traditional approaches, such as Convolutional Neural Networks (CNNs) and Transformers. While CNN-based methods commonly invent elaborate attention mechanisms to grasp the informative features, restrained by the inherent nature of convolution units, they have limited respective fields and cannot capture long-range dependencies. As shown in Fig. 1, the effective receptive field (ERF) [17] of CNN-based model NLSN [18] is limited. This requirement is essential for SR tasks, as a predicted pixel needs prior knowledge from its surrounding region to be super-resolved, especially in wide-range RSI.
Figure 1: The Effective Receptive Field (ERF) [17] comparison for (a) CNN-based method NLSN [18], (b) Transformer-based model RGT [19], and the proposed Mamba-based network FMSR. A wider distribution of dark areas demonstrates larger ERF. Our FMSR effectively obtains the largest ERF, indicating favorable global exploration capability.
Transformer-based methods achieve increased respective fields by leveraging global interaction among all input data through a self-attention (SA) mechanism [20, 21, 22, 23, 24]. Despite achieving superior performance against CNN-based approaches, these methods exhibit quadratic complexity with respect to the token size. In this context, taming Transformer for high-resolution scenarios presents a significant challenge, particularly for large-scale remote sensing images. Although some approaches seek a lightweight SA for global modeling, such as recursive SA [19] and window-based SA [25], they usually come at the expense of global modeling accuracy and require stacking many blocks to establish a global dependency, thus increasing the computational budget. Moreover, the inherent issue of quadratic complexity remains unsolved. Therefore, a natural question arises: can a more efficient yet effective solution be developed to grasp the long-range dependencies across large-scale RSI?
The recent-popular State Space Model (SSM) could be a promising answer to this question. Originating from Kalman filtering [26], SSM primarily employs linear filtering and prediction methods to represent the evolution of the internal state of the system, thus naturally enjoying linear complexity. By integrating SSM with the MLP block of Transformer, the simplified architecture Mamba [27] is achieved, which introduces a selection scanning mechanism in SSM to filter out irrelevant information for long-sequence modeling. Recently, Mamba demonstrated impressive results in various domains, making it a possible replacement for the Transformer model. Nevertheless, some potential problems persist in large-scale earth observation scenarios, making Mamba-based SR more challenging. Firstly, images captured by satellite platforms often lose crucial frequency information for perception, which requires heterogeneous representation for accurate reconstruction. The reason lies in that the original Mamba processes each token equally in the spatial domain, limiting its ability to perceive informative frequency signals. Secondly, there exists spatial diversity among observed objects in RSI, but is barely explored. Although the widely used channel attention mechanism could explore inter-channel relationships, it lacks explicit consideration of spatially-varying contents, resulting in suboptimal pixel-wise representation during local modeling.
To mitigate the aforementioned problem, we first attempt to extend Mamba from the perspective of frequency analysis and propose a frequency-assisted Mamba for RSI SR, termed FMSR. Specifically, instead of solely employing VSSM for long-range modeling, we devise an effective Frequency Selection Module (FSM) to adaptively identify informative frequency cues vital for perception. By incorporating FSM with Mamba, the resulting frequency-assisted mamba block can better utilize the complementary strengths between Mamba and frequency analysis for accurate SR. Considering that the VSSM aggregates image features via patch-wise linear scaling, it inevitably overlooks some pixel-wise localities. To address this, we further develop a Hybrid Gate Module (HGM) to better introduce a local inductive bias. Unlike the commonly used channel attention, HGM allows for selective amplification or attenuation of local features on spatial position, effectively enhancing spatially-varying representation during channel-wise correction learning.
Furthermore, there is inherent misalignment between different level features (global and local). The direct fusion of multi-level features inevitably arises in confused and conflicted representation. To this end, we introduce a learnable adapter to rescale cross-level representation for improved integration. Overall, equipped with the above designs, our FMSR can capture both global and local dual-domain dependencies for RSI SR while maintaining moderate complexity.
Our main contribution can be summarized as follows:
  1. 1)
    We pioneer the first state space model for remote sensing image super-resolution (FMSR), demonstrating the potential of Mamba for efficient yet effective global modeling in large-scale remote sensing scenes.
  2. 2)
    To grasp the frequency dependencies, we devise a Frequency Selection Module (FSM) capable of adaptively selecting the most informative frequency cues through lightweight convolution and activation operations during Fourier transformation.
  3. 3)
    We develop a Hybrid Gate Module (HGM) that benefits the local bias of the CNN operator and spatially-varying coordinates to leverage more faithful localities for reconstruction.
The remainder of this paper is organized as follows. Section II reviews the related knowledge pertinent to our FMSR. In Section III, we provide detailed descriptions of the implementation of our FMSR. Section IV contains extensive experiments conducted on widely used remote sensing benchmarks. Section V concludes our work.
II Related Work
In this section, we first present a comprehensive review of remote sensing image super-resolution. Then, we introduce relevant background knowledge for this study, including state space models and frequency learning.
II-A Remote Sensing Image Super-resolution
RSI SR has witnessed significant advancements with the booming of deep learning [28]. The primary focus of this task lies in extracting prior knowledge from LR images, which can be broadly categorized into three categories: CNN-, Transformer-, and Mamba-based methods.
CNN-based. Drawing inspiration from SRCNN [29], early efforts usually elaborate CNNs with advanced modules, such as residual [30] and dense structure [31] and attention mechanisms [32, 33]. Mei et al. [18] introduced non-local sparse attention to capture global dependencies inherent in LR images. Similarly, Lei et al. [34] extended the non-local mechanism by exploring cross-scale similarity in RSI. While these methods improve the local receptive field nature of CNNs, they suffer from significant computational overhead during non-local exploration, making them less efficient in large-scale remote sensing scenes. Moreover, limited by the local bias of CNN, they cannot capture critical long-range dependencies and reach a plateau in performance.
Transformer-based. The core insight of Transformer lies in the Self-Attention (SA) mechanism, which has demonstrated superior long-range modeling capability and outperformed CNN-based methods. Lei et al. [35] devised a multi-stage Transformer-enhanced network. Recently, Chen et al. [25] proposed to activate more pixels in SR tasks by combining both CNN and self-attention. More Recently, an improved network [36] was proposed by integrating channel attention, improved SA, and anchored stripe attention. However, due to the quadratic complexity of SA regarding token size, they are less efficient in handling high-resolution images, large-scale RSI in particular. To alleviate the computational budget of SA, Chen et al. [19] proposed a recursive SA by recursively aggregating input features for enriched token representation. Nevertheless, efficient SA often sacrifices global modeling capability, and despite efforts to mitigate the quadratic complexity, the inherent problem of SA remains unsolved.
Mamba-based. In light of the success of Mamba, some scholars [37] attempt to introduce Mamba for efficient global modeling with linear complexity. However, to the best of our knowledge, the potential of Mamba in RSI SR remains unexplored. Since the imaging is wide-ranging, the content in RSI exhibits complex and diverse properties. Furthermore, compared to natural images, the texture information of RSI is less prominent, and the vital high-frequency information tends to vanish in deep models.
In summary, there is an urgent need for an efficient yet effective scheme to model the heterogeneous representation and to seek a practical solution to explore the critical high-frequency components. This paper pioneers exploring Mamba’s potential in the RSI SR task and extends Mamba with frequency analysis, providing an effective and efficient paradigm for this challenging issue.
Figure 2: Overview of the proposed FMSR. The Frequency-assisted Mamba Blocks (FMB) are arranged sequentially in Frequency-assisted Mamba Groups (FMG). In FMB, a Frequency Selection Module (FSM) is adopted to assist the learning process of the Vision State Space Module (VSSM) and Hybrid Gate Module (HGM). $\alpha_{l}$ is a learnable adaptor for hybrid adaptive integration in the $l$-th FMB.
II-B State Space Model
Recently, state space models (SSMs) [38] have emerged as a promising approach, demonstrating competitive performance in long-range modeling compared to transformers. The key advantage of SSMs lies in their linear scaling with sequence length, providing a global perspective with linear complexity. Gu et al. [39] pioneered the SSM to tackle long-sequence data modeling, illustrating promising linear scaling properties. Subsequently, they put forward a variant named Mamba [40], which adopts a selective mechanism and efficient network design. Mamba has shown superior performance to transformers in natural language processing tasks. In light of the success of Mamba, it has been introduced in computer vision tasks and demonstrated impressive performance, including object detection [41], image classification [42], and biomedical image segmentation [43]. However, research on Mamba in low-level vision tasks is still in its primary stage, and efforts for RSI SR remain unexplored.
This paper adapts Mamba for the SR task. Unlike previous works that solely replace self-attention with the Vision State Space Model (VSSM) for long-range modeling, we promote the perspective of global exploration in the spatial-frequency dual domain. Compared to spatial-wise modeling, our method is more effective by incorporating latent high-frequency cues for better global representation.
Figure 3: The proposed Hybrid Gate Module (HGM) conceptual illustration. The input feature X is split in the channel dimension and fed through a Channel Attention Block (CAB) and a pixel-wise linear projection layer, respectively. After a Hadamard product operation, a $1\times 1$ convolution generates the output tensor Y.
Figure 4: Three variants of Frequency Selection Module (FSM). Here, we adopt 2D Fast Fourier Transformation (FFT) for frequency learning.
II-C Fourier Transform
The Fast Fourier transform (FFT) can be viewed as a global statistical signal, making it suitable for global information analysis. In light of this, various visual tasks leverage FFT for frequency domain modeling, such as semantic segmentation [44] and image classification [45]. In low-level vision scenes, Mao et al. [46] introduced a residual FFT module capable of capturing comprehensive local high-frequency details for low-light enhancement. Li et al. [47] integrated the Fourier transform into a deep network to mitigate noise amplification during luminance enhancement. Guo et al. [48] proposed a window-based frequency channel attention mechanism. Wang et al. [49] conducted mutual learning of frequency and spatial domains to improve face image SR. Further, some approaches incorporate FFT into the loss function to enhance reconstructed sharp details [50].
However, these methods often prioritize elaborate networks for exploring frequency signals, neglecting frequency contribution analysis. This inevitably amplifies harmful frequencies and increases computational overhead. In this study, we dynamically adjust input frequency-domain features using lightweight activation weights and a convolutional layer to emphasize informative frequencies for accurate SR, which offers more flexibility to modulate selection thresholds.
III Methodology
In this section, we introduce the implementation details of our FMSR. The FMSR comprises a straightforward backbone with convolution layers, Frequency-assisted Mamba Groups (FMG), and a pixel-shuffle layer. We start with an overview of FMSR, and then we dive into its components by explaining: the Frequency-assisted Mamba Block (FMB), Vision State Space Module (VSSM), Hybrid Gate Module (HGM), and Frequency Selection Module (FSM).
III-A Overview of FMSR
As illustrated in Fig. 2, the proposed FMSR consists of three major stages: shallow feature extraction of LR, deep feature acquisition, and reconstruction of HR. Firstly, the given LR input $I_{LR}$ is fed into a $3\times 3$ convolution layer $\phi$ with learnable parameters $\theta$ to generate initial feature $F_{0}$, which can be written as:
${F_{0}}=\phi({I_{LR}},\theta).$ (1)
Subsequently, the $F_{0}$ undergoes deep feature extraction with multiple Frequency-assisted Mamba Groups (FMGs), followed by a $3\times 3$ convolution for feature refinement. This process can be expressed as:
${F_{m}}={\rm FMG}_{m}({\rm FMG}_{m-1}(\cdots{\rm FMG}_{1}({F_{0}}))),$ (2)
where $m$ represents the number of FMG and $F_{m}$ is the output of $m$-th FMG. We incorporate a global skip connection to prepare high-quality features $F_{rec}$ for reconstruction, i.e., ${F_{rec}}={\rm Conv}({F_{m}})+{F_{0}}$. Finally, a $3\times 3$ convolution, a pixel-shuffle layer $\rm PS$, and a terminal convolution are employed to upscale and restore the super-resolved output $I_{SR}$:
${I_{SR}}={\rm Conv}({\rm PS}({\rm Conv}({F_{rec}}),s)),$ (3)
where $s$ means the upscale factor.
III-B Frequency-assisted Mamba Block
The structure of FMB is shown in Fig. 2, from which we can find that FMB serves as a primary component in FMG. In particular, each FMG contains cascaded FMB, a convolution layer, and a residual connection. The FMB is responsible for exploring global and local representations in a frequency-spatial dual domain. The function of $i$-th FMG can be summarized as follows:
${F_{i}}={\rm Conv}(\Psi_{n}(\Psi_{n-1}(\cdots\Psi({F_{i-1}}))))+{F_{i-1}},$ (4)
where $\Psi_{n}$ denotes the $n$-th FMB in each FMG.
Specifically, FMB performs global and local modeling in the frequency-spatial dual domain. Given output of $\Psi_{l-1}$, termed $x_{l-1}$, the l-th FMB $\Psi_{l}$ processes $x_{l-1}$ with three parallel branches to grasp their merits of global representations: 1) A Layer Norm (LN) followed by the 2D Vision State Space Module (VSSM) is developed to capture the spatial-wise long-term information, 2) a Frequency Selection Module (FSM) is equipped to introduce more high-frequency cues in the frequency domain while promoting the capability of VSSM, 3) a learnable scaling factor $\alpha_{l}$ for dynamic feature aggregation. The feature $y$ after global frequency-spatial exploration can be formulated as:
$y={\alpha_{l}}\cdot{x_{l-1}}+{\rm VSSM}({\rm LN}({x_{l-1}}))+{\rm FAB}({x_{l-1}}).$ (5)
After that, $y$ undergoes further local modeling. Similarly, we adopt LN to normalize $y$ and then use a Hybrid Gate Module (HGM) to grasp the spatial locality. Also, FAB is employed to assist the local modeling process with frequency learning. Finally, another scale factor $\alpha_{l+1}$ is used to adaptively integrate the output from local and global representations, which can be expressed as:
$x_{l}={\beta_{l}}\cdot y+{\rm HGM}({\rm LN}(y))+{\rm FAB}(y).$ (6)
TABLE I: Ablation Studies of different components on the proposed FMSR. The best and second-best PSNR performances are highlighted in bold and underlined. Note that we use residual blocks in Model-1 and Model-2 as feed-forward networks.
Components Model-1 (Base) Model-2 Model-3 Model-4 Model-5 Model-6 (FMSR)
Window-based Self-Attention [25]
2D-VSSM (Ours)
MLP Layer [19]
Channel Attention [32]
Hybrid Gate Module (Ours)
Frequency Selection Module (Ours)
PSNR (dB) 27.751 27.846 28.104 28.088 28.122 28.178
III-C Vision State Space Module
Previous efforts often rely on Transformers to explore global dependency, which calculate the long-range response with the self-attention mechanism. Despite achieving favorable performance, they suffer from high complexity, hindering the efficient modeling in large-scale remote sensing images. Inspired by the success of the vision state space module in long-term modeling and aggregation with linear complexity, we first introduce VSSM to the RSI SR task.
In particular, as illustrated in Fig. 2, the normalized feature $x_{N}={\rm LN}(x_{l})$ is expanded along the channel dimension by a linear projection operation $\phi_{1}$ with an expansion factor $\lambda$. Then, a series of operations including a $1\times 1$ Depth-Wise Convolution (DWConv) $f_{1\times 1}$, a SiLU activation $\sigma_{1}$, as well as the 2D-selective scan module (SSM) and LN are sequentially stacked to generate the output of the first branch, denoted $h_{1}$. This branch can be defined as:
${h_{1}}={\rm LN}({\rm SSM}(\sigma_{1}({f_{1\times 1}}(\phi_{1}({x_{N},\lambda}))))).$ (7)
In the second branch, another linear layer $\phi_{2}$ and SiLU function $\sigma_{2}$ are used. The output of this branch can be obtained by:
${h_{2}}={\sigma_{2}}({\phi_{2}}({x_{N}},\lambda)).$ (8)
Finally, to produce the final output $h_{out}$, the output $h_{1}$ and $h_{2}$ are incorporated via Hadamard product, followed by a linear layer $\phi_{3}$. That is:
${h_{out}}={\phi_{3}}({h_{1}}\otimes{h_{2}}).$ (9)
Figure 5: Feature visualization comparisons. The feature maps corresponding to each reference image are the results of the 56th channels in the final FMG.
TABLE II: Ablation studies on different variants of Frequency Selection Module (FSM) as illustrated in Fig. 4. The best and second-best PSNR performances are highlighted in bold and underline, respectively.
Varieties FAB (a) FAB (b) FAB (c)
PSNR (dB) 28.129 28.147 28.178
TABLE III: Ablation studies on Hybrid Adaptive Integration (HAI) of global and local representations. The best and second best PSNR/SSIM performance are highlighted in bold and underline, respectively. We report these results on AID [51].
Method Skip Adaptive $\alpha$Param. PSNR (dB) SSIM
w/o HAI 11.75M 30.80 0.8121
w/ Skip 11.75M 30.57 0.7986
w/ HAI 11.76M 30.93 0.8156
III-D Hybrid Gate Module
Current methods often utilize either MLP layers [25] for feature propagation or incorporate convolution operations, such as attention mechanisms [37], after long-range exploration to introduce critical locality for improved performance. draws inspiration from these approaches but condenses them into a unified hybrid module. To learn a comprehensive local context, the HGM incorporates the spatially-varying properties of RSI by selective amplification or attenuation of local features on the pixel domain while preserving channel-specific features.
As shown in Fig. 3, HGM treats features captured by local convolution as coordinates and then multiplies them with a pixel-wise gating mask of the same size. Specifically, the input feature $\rm X$ is first fed through a $1\times 1$ convolution to expand the channel dimension to $2c$, then split by half in the channel dimension. In the first branch, a $1\times 1$ convolution, a $3\times 3$ depth-wise convolution, and a channel attention block [32] are used to yield the coordinates.
${\rm X}_{coor}={\rm CA}({\rm Dconv}({\rm Conv}({\rm X}_{1}))).$ (10)
In the second branch, we fed the feature through pixel-wise linear projection, somewhat similar to the MLP layer. In contrast to the MLP layer, we only activate the feature at the end of linear projection with the GELU activation function to generate the gate weights:
${\rm M}={\rm GELU}({\rm Linear}({\rm X}_{2})).$ (11)
Finally, the output $\rm Y$ can be obtained by:
${\rm Y}={\rm Conv}({\rm M}\odot{{\rm X}_{coor}}).$ (12)
III-E Frequency Selection Module
To achieve frequency-spatial dual domain representation at global and local levels, we equip VSSM and HGM with frequency exploration. As illustrated in Fig. 4, we devise three variants of frequency selection operations using Fast Fourier transformation (FFT):
(a) Using 2D real FFT and using a $1\times 1$ convolution layer before inverse FFT, which means we do not perform frequency selection:
${\rm Z}={{\cal F}^{-1}}({\rm Conv}({\rm{{\cal F}}}(x))).$ (13)
(b) Inserting ReLU activation between FFT and Inverse FFT to dynamically select the frequency pattern:
${\rm Z}={{\cal F}^{-1}}({\rm ReLU}({\rm{{\cal F}}}(x))).$ (14)
(c) Applying two stacks of $1\times 1$ convolution layer and GELU activation function:
${\rm Z}={{{\cal F}}^{-1}}{\rm Conv}(({\rm GELU}({\rm Conv}({\cal F}(x))))).$ (15)
We finally choose scheme (c) as our FSM as $1\times 1$ convolution lets the network modulate flexible thresholds for frequency selection with lightweight design. Ablation experiments demonstrate the effectiveness of (c) compared to (a) frequency analysis without selection and (b) ReLU-based selection.
Figure 6: Visualization of the feature maps. The proposed Frequency Selection Module (FSM) yields sharp and clear details for reconstruction.
TABLE IV: Ablation studies for the number of FMB. The best and second-best PSNR performances are highlighted in bold and underline, respectively.
Number of FMB PSNR (dB) #Param. GFLOPs
2 27.744 4.55M 58.09G
4 28.026 8.15M 93.18G
6 28.178 11.76M 128.27G
8 28.181 15.37M 163.36G
TABLE V: Ablation studies for the number of FMG. The best and second-best PSNR performances are highlighted in bold and underline, respectively.
Group 2 4 6 8
#Param. 4.21M 7.99M 11.76M 15.54M
GFLOPs 52.66G 90.46G 128.27G 166.07G
PSNR (dB) 27.691 27.985 27.178 28.154
Figure 7: Ablation studies of memory consumption, inference times, parameters, and PSNR performance on DOTA [52]. Note that the inference times are calculated on 100 images.
TABLE VI: Model efficient comparison. The FLOPs results are calculated with an input tensor of size $1\times 3\times 128\times 128$. The inference times are tested on 100 random images.
Metrics EDSR [53] RCAN [32] HAN [33] HSENet [34] NLSN [18] TransENet [35] HAT-L [25] RGT [19] MambaIR [37] FMSR
Parameters (M) 43.09 15.59 16.07 21.7 44.15 37.46 40.32 10.19 16.72 11.76
FLOPs (G) 823.34 261.01 268.89 306.31 840.79 87.85 672.15 40.12 280.82 128.27
Memory (MB) 202 62 82 108 206 158 164 72 92 46
Inference Times (ms) 79.188 136.71 149.24 59.95 125.42 91.09 214.56 191.48 60.99 100.88
TABLE VII: Quantitative comparison on AID [51], DOTA [52], and DIOR [54] test set in terms of PSNR, SSIM, and LPIPS, where the 1st, 2nd, and 3rd best performance are highlighted in red, blue, and green, respectively. FMSR++ means the self-embedding results of our FMSR.
Methods Venue AID [51] DOTA [52] DIOR [54] Average
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Bicubic - 28.86 0.7382 0.4803 31.16 0.7947 0.4043 28.57 0.7432 0.4678 29.53 0.7587 0.4508
EDSR [53] CVPRW’2017 30.65 0.8086 0.3068 33.64 0.8648 0.2616 30.63 0.8116 0.3020 31.64 0.8283 0.2901
RDN [31] CVPR’2018 30.74 0.8112 0.3157 33.60 0.8670 0.2421 30.78 0.8147 0.3093 31.71 0.8310 0.2890
RCAN [32] ECCV’2018 30.82 0.8121 0.3112 33.86 0.8680 0.2384 30.85 0.8159 0.3048 31.84 0.8320 0.2848
HAN [33] ECCV’2020 30.80 0.8122 0.3079 33.84 0.8682 0.2363 30.84 0.8163 0.3026 31.83 0.8322 0.2823
HSENet [34] TGRS’2022 30.72 0.8108 0.3124 33.85 0.8667 0.2412 30.77 0.8143 0.3070 31.78 0.8306 0.2869
NLSN [18] CVPR’2021 30.81 0.8126 0.3044 33.86 0.8682 0.2349 30.82 0.8156 0.3004 31.83 0.8321 0.2799
HAUNet [55] TGRS’2023 30.88 0.8132 0.3111 33.94 0.8687 0.2403 30.87 0.8160 0.3074 31.90 0.8326 0.2863
TransENet [35] TGRS’2022 30.80 0.8109 0.3136 33.75 0.8675 0.2418 30.85 0.8148 0.3089 31.80 0.8311 0.2881
HAT-L [25] CVPR’2023 30.81 0.8124 0.3078 33.99 0.8684 0.2392 30.87 0.8161 0.3062 31.89 0.8323 0.2844
GRL-L [36] CVPR’2023 30.86 0.8127 0.3085 33.86 0.8710 0.2372 30.90 0.8177 0.3047 31.87 0.8338 0.2835
RGT [19] ICLR’2024 30.91 0.8159 0.3023 33.92 0.8709 0.2337 30.91 0.8182 0.2992 31.91 0.8350 0.2784
MambaIR [37] Arxiv’2024 30.85 0.8130 0.3098 33.86 0.8691 0.2388 30.89 0.8167 0.3060 31.87 0.8329 0.2849
FMSR (Ours) - 30.93 0.8156 0.3001 34.04 0.8710 0.2325 30.97 0.8187 0.2983 31.98 0.8351 0.2770
FMSR++ (Ours) - 31.07 0.8185 0.3067 34.27 0.8735 0.2363 31.13 0.8219 0.3039 32.16 0.8380 0.2823
Figure 8: The visualization of Local Attribution Maps (LAM) [56]. A wider range of LAM illustrates more pixels are involved in reconstruction. A higher diffusion index demonstrates better global activation capability.
TABLE VIII: Quantitative comparison with SOTA CNN-, Transformer-, and Mamba-based SR methods across 30 scene categories on AID [51], where the best and second best PSNR/SSIM performance are highlighted in red and blue, respectively.
Scene Categoties Bicubic EDSR [53] HSENet [34] NLSN [18] HAT-L [25] MambaIR [37] FMSR (Ours)
[…]
Airport 27.83 0.7554 29.93 0.8282 30.08 0.8303 30.16 0.8322 30.15 0.8319 30.21 0.8327 30.26 0.8346
Bare Land 35.60 0.8564 36.94 0.8837 36.79 0.8841 37.00 0.8845 36.88 0.8841 36.95 0.8843 37.02 0.8847
Baseball Field 31.00 0.8305 33.05 0.8765 33.15 0.8774 33.24 0.8787 33.25 0.8789 33.32 0.8792 33.43 0.8803
Beach 32.90 0.8446 34.18 0.8727 34.14 0.8746 34.31 0.8749 34.34 0.8756 34.36 0.8754 34.38 0.8759
Bridge 30.22 0.8283 32.93 0.8800 33.06 0.8809 33.12 0.8818 33.04 0.8809 33.14 0.8821 33.28 0.8836
Center 26.51 0.6944 28.77 0.7921 28.83 0.7937 28.95 0.7971 28.92 0.7956 28.99 0.7966 29.09 0.8010
Church 24.29 0.6333 26.30 0.7469 26.47 0.7507 26.51 0.7528 26.56 0.7532 26.61 0.7543 26.64 0.7567
Commercial 27.33 0.7174 29.01 0.7940 29.21 0.7989 29.21 0.7996 29.21 0.8007 29.29 0.8012 29.34 0.8039
Dense Residential 22.93 0.5671 24.38 0.6839 24.60 0.6912 24.60 0.6936 24.67 0.6936 24.74 0.6965 24.78 0.7019
Desert 39.26 0.9100 40.20 0.9268 39.57 0.9271 40.27 0.9278 40.37 0.9278 40.09 0.9270 40.29 0.9275
Farmland 33.10 0.8226 35.00 0.8683 35.02 0.8692 35.10 0.8699 35.03 0.8691 35.09 0.8696 35.18 0.8711
Forest 28.79 0.6605 29.85 0.7315 30.00 0.7363 29.98 0.7369 30.01 0.7363 30.00 0.7370 30.04 0.7391
Industrial 26.77 0.6952 28.88 0.7931 28.98 0.7956 29.04 0.7982 29.04 0.7980 29.09 0.7988 29.17 0.8027
Meadow 33.86 0.7483 34.63 0.7804 34.55 0.7804 34.69 0.7821 34.70 0.7815 34.64 0.7811 34.74 0.7827
Medium Residential 26.36 0.6335 28.34 0.7365 28.45 0.7390 28.49 0.7418 28.46 0.7408 28.59 0.7435 28.65 0.7472
Mountain 29.51 0.7349 30.63 0.7885 30.72 0.7907 30.74 0.7916 30.78 0.7923 30.76 0.7918 30.78 0.7928
Park 29.06 0.7530 30.54 0.8130 30.71 0.8167 30.71 0.8177 30.71 0.8189 30.76 0.8185 30.80 0.8204
Parking 24.24 0.7060 27.25 0.8317 27.32 0.8341 27.57 0.8408 27.56 0.8405 27.65 0.8414 28.02 0.8524
Playground 32.64 0.8450 35.37 0.8943 35.46 0.8952 35.58 0.8967 35.49 0.8959 35.54 0.8961 35.72 0.8978
Pond 30.70 0.8167 32.11 0.8542 32.17 0.8549 32.22 0.8559 32.18 0.8555 32.24 0.8559 32.23 0.8560
Port 26.67 0.7986 28.50 0.8596 28.71 0.8623 28.71 0.8631 28.81 0.8638 28.83 0.8644 28.92 0.8668
Railway Station 26.78 0.6793 28.72 0.7738 28.84 0.7762 28.89 0.7783 28.88 0.7780 28.99 0.7802 29.04 0.7830
Resort 26.79 0.7029 28.52 0.7799 28.64 0.7825 28.68 0.7845 28.71 0.7849 28.76 0.7857 28.82 0.7886
River 30.37 0.7402 31.55 0.7891 31.61 0.7904 31.64 0.7914 31.63 0.7909 31.64 0.7912 31.66 0.7918
School 27.41 0.7237 29.36 0.8044 29.51 0.8074 29.55 0.8097 29.54 0.8104 29.62 0.8113 29.70 0.8143
Sparse Residential 26.66 0.6006 27.71 0.6728 27.84 0.6754 27.84 0.6767 27.88 0.6759 27.88 0.6768 27.91 0.6789
Square 28.55 0.7391 30.84 0.8200 30.94 0.8223 31.04 0.8244 31.00 0.8251 31.05 0.8247 31.16 0.8276
Stadium 27.16 0.7547 29.63 0.8387 29.68 0.8391 29.79 0.8422 29.77 0.8422 29.83 0.8421 29.97 0.8460
Storage Tanks 25.65 0.6793 27.44 0.7664 27.58 0.7688 27.61 0.7709 27.60 0.7698 27.64 0.7709 27.71 0.7739
Viaduct 26.97 0.6755 28.99 0.7757 29.08 0.7772 29.17 0.7813 29.11 0.7794 29.19 0.7810 29.26 0.7840
Average 28.86 0.7382 30.65 0.8086 30.72 0.8108 30.81 0.8126 30.81 0.8124 30.85 0.8130 30.93 0.8156
Figure 9: Visual comparisons of our FMSR with CNN-, Transformer-, and Mamba-based methods on AID [51] with scale ×$\times 4$. Zoom in for better observation.
Figure 10: Visual comparisons of our FMSR with CNN-, Transformer-, and Mamba-based methods on DOTA [52] with scale ×$\times 4$. Zoom in for better observation.
Figure 11: Visual comparisons of our FMSR with CNN-, Transformer-, and Mamba-based methods on DIOR [54] with scale ×$\times 4$. Zoom in for better observation.
IV Experiment
IV-A Datasets
In this paper, we report the results of the SR performance on three RSI benchmarks, including AID [51], DOTA [52], and DIOR [54]. In particular, AID is used to form the training and test simultaneously. We randomly select 3000 and 900 images from AID to form the training and test set, with an image size of $640\times 640$. Note that the training and test parts of AID are non-overlapping. In DOTA and DIOR, 900 and 1000 images are randomly extracted for model evaluation, respectively. Both of them are with a size of $512\times 512$.
IV-B Implementation Details
Model Details. This paper focuses on ×$\times 4$ SR. Our FMSR is constructed by 6 FMGs for deep feature exploration, with each FMG consisting of 6 FMBs, i.e., $m=n=6$. Empirically, we set the internal channel dimension to $c=96$. ll convolutional kernel sizes are set to $3\times 3$, except for those in the Hybrid Gate Module (HGM) and Frequency Selection Module (FSM), which utilize $1\times 1$ kernels for increased efficiency. The expansion rate in the linear projection layer is set to $\lambda=2$.
Training Details. During the training procedure, all the SR methods were retrained on the AID training set using L1 loss and ADAM algorithm with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. To train our FMSR, we randomly select 4 image patches with the size of $64\times 64$ in each mini-batch. The learning rate is initialized to $1\times 10^{-4}$ and halved every 200 epochs until training stops at 500 epochs. All SR models were implemented in the PyTorch framework and trained on a single NVIDIA RTX 3090 GPU with 24 GB memory and a 3.40 GHz AMD Ryzen 5700X CPU.
IV-C Evaluation Metrics
Two classical full-reference indicators are used to evaluate the SR performance, i.e., Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [57]. Moreover, the LPIPS metric [58] is employed to analyze the perceptual quality of super-resolved results. Note that PSNR and SSIM are calculated on the luminance channel (Y) of YCbCr space.
IV-D Ablation Study
In this section, we discuss the proposed FMSR in depth by investigating the effect of its major components and their variants. All these models are trained on AID with scale factor ×$\times 4$. Following prior work [59], we employ a small-scale dataset, AID-tiny, consisting of 30 randomly selected images from AID, for efficient evaluation unless otherwise specified. The baseline model is derived by excluding FSM and substituting VVSM and HGM with the standard window-based self-attention and 5 residual blocks, respectively.
IV-D1 Effect of Key Modules
(a) Effect of VSSM. Table I reports that the Model-1 (baseline) obtains 27.751 dB. Model-2 exhibits a performance gain of 0.089 dB over the baseline model, which demonstrates the effectiveness of VSSM in global modeling. In addition, in Fig. 5, we visualize the intermediate feature maps of the baseline model and our FMSR. The features are obtained by visualizing the 56th channels at the end of FMG. With VSSM, the results of our FMSR are more prominent across the entire feature maps, highlighting the favorable global modeling capability.
(b) Effect of HGM. Employing both VSSM and the MLP layer, Model-3 demonstrates a favorable performance improvement of 0.258 dB. However, when equipped with the channel attention mechanism to introduce more locality, Model-4 performs similarly to Model-2 (28.088 dB vs. 28.104 dB). This suggests that neither MLP nor channel attention is less effective in boosting the performance of Mamba. In this context, further improving the reconstruction performance becomes very challenging. Our HGM achieves a performance gain of 0.034 dB. Thus, to benefit from both global and local inductive bias, we combine VSSM and HGM in our FMSR.
(c) Effect of FSM. By comparing the results of FMSR and Model-5, we can evaluate the performance of the proposed FSM. In this case, our FSM can bring an improvement of 0.056 dB. As reported in Table II, we discuss some variants of FSM shown in Fig. 4. If we do not perform frequency selection, FSM(a) produces the worst PSNR performance. By inserting a simple ReLU activation, FSM(b) could adaptively eliminate noisy frequencies and achieve a gain of 0.018 dB, demonstrating the effectiveness of the selection mechanism. If we adopt $1\times 1$ convolution and GELU for selection, FMSR allows for a more flexible threshold for frequency selection, thus obtaining the best performance. Moreover, visual comparisons between FMSR and Model-5 are shown in Fig. 6. In Fig. 6, our FMSR generates superior textures in the high-frequency of the bridge. In contrast, without performing frequency selection, the intermediate features are blurred at the boundary and unclear at the edge. These results confirm that our FSM has the ability to explore more critical high-frequency cues for better SR performance.
IV-D2 Effect of HAI
We show the influence of HAI in Table III, where we conduct three ablation analyses: 1) without HAI, 2) with residual connection (Skip), and with HAI. We observe that without HAI, the performance of FMSR drops by 0.13 dB. Additionally, comparing the model with Skip, our FMSR achieves a significant improvement of 0.36 dB. This may be because of the misalignment between global and local representations, which means simply adding them may not elaborate enough to integrate these different levels of knowledge, thus generating suboptimal performance. Benefiting from the adaptive scaling factor $\alpha$, our FMSR could dynamically adjust the features at the global and local ranges, thus generating enriched feature integration.
IV-D3 Model Efficiency
The parameters, FLOPs, and SR performance of state-of-the-art (SOTA) methods are reported in Table VII. Intuitively, we plot the performance versus parameters in Fig. 7(c), where we observe that FMSR strikes a favorable trade-off between performance and parameters. Here, we further investigate the relationship between network structure and model complexity. Moreover, more intuitive metrics are involved to analyze the model efficiency, such as inference times and memory consumption.
(a) Number of FMB and FMG. We study the inference of the number of FMB and FMG in Table IV and Table V, respectively. As we can see in Table IV, using more FMB leads to the consistently increasing PSNR performance from 27.744 to 28.181 dB, demonstrating the effectiveness of FMB. Meanwhile, the parameters grow dramatically from 4.55 to 15.37M. To strike a favorable balance between performance and model size, we finally picked 6 FMB in our FMSR. Regarding the number of FMG, as shown in Table V, the increasing number of groups leads to higher PSNR values. Nevertheless, the performance saturates at group 8, which may be caused by over-fitting. Ultimately, we insert 6 FMG in the FMSR for deep feature learning.
(d) Memory Consumption. The max CUDA memory consumption of our FMSR and SOTA models are shown in Fig. 7(a), from which we can see that our consumes the least memory consumption while outperforming other methods. Specifically, compared to competitive NLSN that adopts non-local attention for global modeling, we find the performance of FMSR is 0.18 dB higher, but also reduces the memory by 160 MB. Furthermore, FMSR achieves better performance against impressive HAT-L with only 28% of its memory. This indicates that FMSR has stronger global modeling capability and greatly surpasses Transformer-based models in model efficiency.
(c) Inference Times. Regarding the inference times, as shown in Fig. 7(b), our FMSR shows a trade-off with other methods. For instance, FMSR achieves the best PSNR performance compared to the advanced Transformer-based model RGT by 0.12 dB, but at the cost of increased inference time usage of about 90.6 ms, measured by the test times of 100 images. These results provide intuitive evidence that our FMSR is very efficient in large-scale remote sensing image SR tasks.
IV-E Comparisons With State-of-the-Art
IV-E1 Comparative Methods
To evaluate the SR performance of our FMSR against SOTA methods on remote sensing images, advanced CNN-, Transformer-, and Mamba-based models are involved for comprehensive comparison, including EDSR [53], RDN [31], RCAN [32], HAN [33], NLSN [18], HSENet [34], TransENet [35], HAT-L [25], GRL-L [36], RGT [19], and MambaIR [37]. We also report the self-ensemble results of our FMSR, dubbed FMSR++.
IV-E2 Quantitative Evaluations
The quantitative results on the AID, DOTA, and DIOR datasets are shown in Table. VII. Our FMSR++ and FMSR achieve the highest and second-highest average performance on all metrics, demonstrating their superior SR performance across various remote sensing benchmarks. Particularly on the AID dataset, FMSR generates a substantial gain of 0.12 dB PSNR over HAT-L with only 29% parameters and 19% FLOPs. Compared to GRL-L, FMSR obtains 0.18 dB higher PSNR on the DOTA dataset with 60% lower complexity. Our FMSR exhibits 0.3001 LPIPS on AID, which is much higher than other comparative approaches. Moreover, recent Mamba-based methods, like MambaIR, achieved marginal improvements against CNN-based models. For instance, compared to NLSN, which explores global dependency with non-local attention, MambaIR only leads NLSN by 0.04 dB on AID. This underscores the effectiveness of VSSM in exploring non-local dependency. However, it also indicates that spatial-wise global modeling reaches a plateau in complex remote sensing images. In contrast to simply employing VSSM for global modeling, our FMSR seeks a more practical solution in the frequency domain, thus achieving significant improvement over MambaIR.
In addition, to validate the generalization capability of SR models on diverse remote sensing senses, we further report the PSNR and SSIM results on AID across 30 scene categories. The results are shown in Table VIII. As we can see, FMSR demonstrates stronger generalization capability against SOTA methods and obtains the best performance on almost all remote sensing scenarios, receiving 0.18 dB PSNR and 0.039 SSIM over MambaIR on the ”Industrial” scene. Furthermore, in the ”Stadium” category, our FMSR significantly outperforms HAT-L by a large margin. These notable improvements obtained by FMSR align with our motivation, which aims to introduce Mamba for efficient yet favorable global modeling in large-scale remote sensing images.
IV-E3 Qualitative Results
Visual comparisons on AID, DOTA, and DIOR with a scale factor of ×$\times 4$ are shown in Fig. 9, Fig. 10, and Fig. 11. From these visualizations, we observe that our FMSR recovers sharp edges with richer textures, especially capturing critical high-frequency details in these remote sensing images. For example, comparing the restored ”playground_270” in Fig. 9, we can see that the recent competitive Transformer-based method HAT-L and Mamba-based model MambaIR struggle to reconstruct the HR lines on the runway. Moreover, as illustrated in ”img_165” and ”img_658” of Fig. 10, only the proposed FMSR can recover severely damaged lines on the ground, while other SR models fail to produce accurate distribution on the details. These visual comparisons further demonstrate the effectiveness of our FMSR in capturing and reconstructing fine details in RSI.
When comparing the visual results of the top image from the DIOR dataset in Fig. 11, where large-scale and global information exists, we observe that all CNN-, Transformer-, and Mamba-based SR networks struggle to handle this issue, resulting in suboptimal results and losing some high-frequency textures. In contrast, benefiting from the global modeling capability of VSSM as well as the spatial-frequency dual-domain exploration, our FMSR produces results that are visually close to the ground truth and successfully generate multi-frequency lines. For example, MambaIR without frequency selection cannot recover accurate details, while HAT-L produces severe distortion. Our FMSR still maintains favorable visual quality with more high-frequency contextual information. Similarly, the visual comparisons in the bottom image of Fig. 11 provide additional evidence of the superiority of our FMSR.
In addition, the LAM comparisons shown in Fig. 8 demonstrate that FMSR can exploit more pixels during the SR process thanks to the favorable long-range modeling capability of VSSM. By comparing FMSR with HAT-L, FMSR surpasses HAT-L by 8.637 in terms of the diffusion index, indicating the strong wide-range pixel utilization capability due to our spatial-frequency dual-domain representation.
V Conclusion
In this study, we first introduce the state space model for remote sensing image super-resolution. FMSR effectively models global dependencies in large-scale remote sensing images while enjoying the linear complexity. Specifically, we develop an efficient yet effective frequency selection module (FSM) to incorporate more relevant frequencies during spatial-frequency dual-domain learning. Meanwhile, channel-wise attention metrics are linearly scaled to enrich the spatially varying representation. Furthermore, to combine global and local representations, we employ a learnable adaptor to adaptively adjust features across different levels. Extensive quantitative and qualitative experiments across AID, DOTA, and DIOR benchmarks demonstrate the superior performance of our FMSR in remote sensing image super-resolution tasks compared to state-of-the-art CNN-, Transformer-, and Mamba-based SR models.